Find n such as Tn approximates \int_{1}^{2}\frac{1}{x+1}dx with a 15 digit precision. What irrational number is approximating?
Answer #1

I du +1 2 こ Query E Subatitute t= 1+1 dt = du. I felt = logns - lagg-loga log (3) -0.4054651081081 644, Hence TM= 0.405465108